Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Three points are chosen independently an at random on the ci

Expert replies
by BTGmoderatorDC » Thu Feb 14, 2019 3:45 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

Three points are chosen independently an at random on the circumference of a circle with radius r. What is the approximate probability that none of the three points lies more than a straight-line distance of r away from any other of the three points?

(A) 1/9
(B) 1/12
(C) 1/18
(D) 1/24
(E) 1/27

OA B

Source: Manhattan Prep
Join the discussion
Source: — Problem Solving |

BTGmoderatorDC wrote:Three points are chosen independently an at random on the circumference of a circle with radius r. What is the approximate probability that none of the three points lies more than a straight-line distance of r away from any other of the three points?

(A) 1/9
(B) 1/12
(C) 1/18
(D) 1/24
(E) 1/27
Source: Manhattan Prep
Image

Very nice problem!

FOCUS : the probability of having all three points chosen (say blue, red and green, in that order) "sufficiently close".

1. The first point (blue) may be considered (without loss of generality) placed at the top of the circle (first figure).

2. The second point (red) must be placed in the arc of the circle shown in red, somewhere between the two nearer vertices of the top vertex of the regular hexagon (first figure).

Reason: a regular hexagon is composed of 6 equilateral triangles, in this case each side with length r (the radius of the circle that circumscribes the hexagon).

Conclusion: there is a 120/360 = 1/3 probability of choosing the second point favorably.

3. What about the third (green) point? Let´s start considering the two extremal scenarios:

> Extremal Case 1 (where the blue and red points coincide): in this case, we have again 1/3 of probability of choosing the third point in a place such that we have all three points "sufficiently close".

> Extremal Case 2 (where the blue and red points are most apart): in this case, we have 60/360 = 1/6 of probability of choosing the third point in a place such that we have all three points "sufficiently close".

> What about an "Intermediate Case"? If we move the red point away from the blue point, from the Extremal Case 1 until the Extremal Case 2, we have ALL (the infinite continuum possibilities) probabilities from 1/3 to 1/6 possible and, by symmetry, equiprobable!

Conclusion: we must agree that the probability of choosing the third point favorably must be the AVERAGE of 1/3 and 1/6, that is, (1/3 + 1/6)/2 = 1/4.


Finally, the answer we are looking for:

? = P(choosing the second point wisely) * P(choosing the third point wisely given that the second one was chosen wisely) = 1/3 * 1/4 = 1/12.


The correct answer is (B).


We follow the notations and rationale taught in the GMATH method.

Regards,
Fabio.
Fabio Skilnik :: GMATH method creator ( Math for the GMAT)
English-speakers :: https://www.gmath.net
Portuguese-speakers :: https://www.gmath.com.br
Join the discussion